OpenStudy (anonymous):

Help with Math Question!! Will give medals!! If an object is dropped from a height of 200 feet, the function h(t)=-16t^2+200 gives the height of the object after t seconds. Approximately, when will the object hit the ground.

OpenStudy (hoblos):

when the object hit the ground, its height will be equal to 0 thus h(t) = 0 so we have to solve -16t^2 + 200 = 0 do you know how to solve it ?

OpenStudy (anonymous):

Not exactly, I do know that we need to square 16 and 200 but thats it.

OpenStudy (hoblos):

you have two methods here you either use the quadratic formula or you can use the identity a^2 - b^2 = (a-b)(a+b)

OpenStudy (hoblos):

you want to use the second method ?

OpenStudy (anonymous):

I know the quadratic function very well.

OpenStudy (hoblos):

the quadratic equation has the form ax^2 + bx + c so the number multiplying x^2 is a so the number multiplying x is b so the last number is c

OpenStudy (hoblos):

what is multiplying x^2 here ?

OpenStudy (anonymous):

-16

OpenStudy (hoblos):

correct what's multiplying x ?

OpenStudy (anonymous):

200?

OpenStudy (hoblos):

is there any x in the equation ?

OpenStudy (anonymous):

-16t^2+200, no

OpenStudy (hoblos):

when you don't see x in the equation, it can be written like this -16x^2 +0x + 200 thus b=0

OpenStudy (anonymous):

Ohhh right, I understand.

OpenStudy (hoblos):

so can you apply the quadratic formula now ?

OpenStudy (anonymous):

184.00 seconds?

OpenStudy (hoblos):

can you show me your calculation ?

OpenStudy (anonymous):

-16(1)^2+0(1)+200?

OpenStudy (hoblos):

what's this ?

OpenStudy (hoblos):

do you know the quadratic formula ?

OpenStudy (hoblos):

\[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

Yes I know that

OpenStudy (hoblos):

apply it

OpenStudy (anonymous):

Our B is 0 correct?

OpenStudy (hoblos):

yes

OpenStudy (anonymous):

so x= -0+-sq0^2-4*-16*200/2*-16

OpenStudy (anonymous):

3.54 seconds

OpenStudy (hoblos):

correct

OpenStudy (anonymous):

Your a good person for sticking with me! :)

OpenStudy (anonymous):

Thanks -_-

OpenStudy (hoblos):

glad to help :)